# SlowFast-LLaVA-1.5: A Family of Token-Efficient Video Large Language Models for Long-Form Video Understanding

### v1

@@ Abstract
We introduce SlowFast-LLaVA-1.5 (abbreviated as SF-LLaVA-1.5), a family of video large language models (LLMs) offering a token-efficient solution for long-form video understanding. This model family employs the two-stream SlowFast mechanism, enabling efficient modeling of long-range temporal context to meet the demand for lightweight, mobile-friendly Video LLMs. We provide models ranging from 1B to 7B parameters, optimized through a streamlined training pipeline and a high-quality data mixture composed of publicly available datasets. Experimental results demonstrate that SF-LLaVA-1.5 achieves competitive performance on a wide range of video and image benchmarks, with robust results across all model sizes. Notably, SF-LLaVA-1.5 achieves state-of-the-art results in long-form video understanding (e.g., LongVideoBench and MLVU) and excels at small scales (1B and 3B) across various video benchmarks.
@@ 1 Introduction
Video large language models (LLMs) (Maaz et al., 2024b; Lin et al., 2023a; Xu et al., 2024a) integrates video perception into pre-trained LLMs to process videos and generate responses to user commands. Although significant progress has been made, notable limitations remain in existing Video LLMs. First, they enhance perception and reasoning over long video sequences by leveraging the LLM’s increasing context length and handling massive input frames (Shen et al., 2024; Chen et al., 2024d; Zhang et al., 2024c). However, the potential for transferring this capability to highly efficient models is underexplored. Second, achieving optimal performance typically requires internal datasets and a complex training lifecycle, with selective parameters frozen at each stage (Li et al., 2024a; Zhang et al., 2025a). These intricate designs result in high computational costs and reproducibility challenges. Third, many Video LLMs (Zohar et al., 2024; Li et al., 2024d) are optimized exclusively for video tasks, limiting their effectiveness as joint models for image understanding tasks.
Building upon the success of SlowFast-LLaVA (Xu et al., 2024b), we introduce SlowFast-LLaVA-1.5, a family of Video LLMs designed for long-form video understanding, focusing on the most efficient model scales (1B and 3B). Our model family is both effective and token-efficient in modeling long-range temporal context. This is achieved by employing the SlowFast mechanism, which balances the trade-off between processing more input frames that significantly increases the token count and computational cost, and reducing tokens per frame that inevitably loses fine-grained details. Specifically, the Slow pathway captures detailed spatial features at a low frame rate, while the Fast pathway operates at a high frame rate with fewer tokens per frame to focus on motion cues. The success of our model also relies on a streamlined training pipeline and a high-quality data mixture. Notably, our model training consists of only two stages. The first stage is supervised fine-tuning on image-only data, providing a good foundation for general knowledge and reasoning. The second stage conducts video-image joint training to learn spatial and temporal features for video understanding while maintaining strong image understanding performance. To ensure seamless reproducibility, all pre-trained weights and training datasets used in this work are publicly accessible.
@@
Our main contributions are as follows. First, we introduce SlowFast-LLaVA-1.5, a family of Video LLMs ranging from 1B to 7B parameters. We demonstrate the effectiveness of incorporating the SlowFast mechanism into a supervised fine-tuning framework, modeling long-range context while maintaining high computational efficiency. Second, our model family requires only two training stages and uses only publicly available datasets, making it fully reproducible. Third, our models achieve the state-of-the-art performance on long-form video understanding. Especially, our smaller models (1B and 3B) clearly outperform comparable Video LLMs across video benchmarks.
@@ 2 Related Work
Image Large Language Models have gained widespread attention (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023; Chen et al., 2024e; Bai et al., 2025). Significant progress across multiple fronts includes: i) enhancing data quantity and quality during pre-training (McKinzie et al., 2024; Liu et al., 2024a; Lin et al., 2023b; Li et al., 2024c) and supervised fine-tuning (SFT) (Zhang et al., 2025b; Deitke et al., 2024; Chen et al., 2024a; Wang et al., 2023; Tong et al., 2025); ii) accommodating images of various high resolutions (Lin et al., 2023c; Zhang et al., 2024b; Wang et al., 2024b); iii) improving architecture designs, including different visual encoders (Zhai et al., 2023; Tong et al., 2024; Shi et al., 2024) and vision-language connectors (Li et al., 2023a; Cha et al., 2024); and iv) conducting comprehensive studies for easy deployment (Team et al., 2023; Marafioti et al., 2025). These rapid advancements also establish a strong foundation for related areas such as video understanding (Maaz et al., 2024b; Lin et al., 2023a), referring & grounding (You et al., 2023; 2024), and visual agents (Durante et al., 2024; Yang et al., 2025).
Video Large Language Models have become an active research area (Li et al., 2023b; Song et al., 2024; Chen et al., 2024b; Zhang et al., 2024e; Zohar et al., 2024). Early work develops Video LLMs as specialist models (Zhang et al., 2023; Cheng et al., 2024; Xu et al., 2024a; Ryoo et al., 2024), achieving strong performance on video tasks but with some trade-offs in image understanding. Training-free Video LLMs (Kim et al., 2024; Xu et al., 2024b) offer an efficient alternative by leveraging Image LLMs without additional fine-tuning on video data, enabling flexible deployment across various applications. Recent models (Zhang et al., 2024g; Liu et al., 2025; Zhang et al., 2025a) are jointly trained on image and video datasets, enabling strong performance in both modalities. Long-form video understanding (Zhou et al., 2024; Wu et al., 2025) also gained a lot of attentions, addressing hour-long videos (Chen et al., 2024d; Li et al., 2024d) or live streams (Qian et al., 2024; Zhang et al., 2024a) while optimizing the token efficiency (Lee et al., 2024b). The proposed SlowFast-LLaVA-1.5 is a family of Video LLMs designed for long-form video understanding. It enhances SlowFast-LLaVA (Xu et al., 2024b) by integrating the SlowFast design into a unified video-image training pipeline, achieving state-of-the-art performance with efficient token consumption.
@@
We provide a detailed explanation of SlowFast-LLaVA-1.5 (abbreviated as SF-LLaVA-1.5), which incorporates the SlowFast video projector into a LLaVA-style architecture, improving long-range temporal modeling while optimizing token efficiency. In contrast to its training-free pioneer (Xu et al., 2024b), this paper i) systematically investigates different instantiations based on the generic SlowFast idea (Sec. 3.1), ii) designs a compact yet effective training pipeline (Sec. 3.2), and iii) introduces tailored data mixtures using only publicly available datasets for each training stage (Sec. 3.3).
@@ 3.1 Model Architecture
As shown in Fig. 1, the architecture of SF-LLaVA-1.5 follows the core design principle of SF-LLaVA (Xu et al., 2024b). It takes a video/image $\mathbf{V}$ and a question $\mathbf{Q}$ as inputs and responds with a textual answer $\mathbf{A}$. For video input, we sample $N$ frames, $\mathbf{I}=\{I_{1},I_{2},...,I_{N}\}$, at a fixed frame rate without special frame assembling ($N$ equals $1$ for image input). After that, a visual encoder (e.g., OryxViT (Liu et al., 2025)) is used to extract frame-level features $\mathbf{F}_{v}\in\mathbb{R}^{N\times H\times W}$ from the inputs independently, keeping their original aspect ratio. The video and image feature tokens are then fed into different projectors, with video using the two-stream SlowFast projector and image using a two-layer MLP.
@@ 3.2 Training Pipeline
Stage I (image understanding) conducts SFT with images to provide a good warmup status for video understanding. For simplicity and efficiency, we do not use any extra pretraining stages (Li et al., 2024a) or image splitting strategy (Lin et al., 2023c), although these are proven to be effective for boosting text-rich results. Instead, we use simple native resolution inputs following Oryx (Liu et al., 2025), where, for each image $I_{i}\in\mathbb{R}^{H_{i}\times W_{i}}$, we have a low resolution $I_{i}^{low}$ and a high resolution $I_{i}^{high}$ inputs. The low-resolution image is obtained by simply resizing the original image to a base resolution, as in $I_{i}^{low}=\textnormal{resize}(I_{i},H_{i}^{base}\times H_{i}^{base})$. For $I_{i}^{high}$, we keep its original aspect ratio and resize it to $H_{i}^{high}\times W_{i}^{high}$, as in Eq. 1 and 2,
@@ 4.1 Implementation Details
  • Training Stage II continues training based on SF-LLaVA-1.5-Image by combining our video and image data mixture. For image, the high-resolution image is obtained in the same way as Stage I, except that we increase $\theta^{I^{max}}$ to $1536^{2}$. For video, we follow prior work (Zohar et al., 2024) and sample frames in 1 FPS. We set the max frame number to 128 and uniformly sample the frames if the number exceeds this upper bound. For each video frame, we set $\theta^{V^{min}}=288^{2}$ and $\theta^{V^{max}}=480^{2}$. The maximum context length is set to 16K. The models trained by this stage are named as SF-LLaVA-1.5.
@@ 4.3 Image Understanding Results
Second, SF-LLaVA-1.5 outperforms strong baselines at the 7B scale, except for InternVL2.5 and Qwen2-VL. Using MM1.5-7B as an example, SF-LLaVA-1.5 achieves better results across benchmarks (e.g., +7.2% on MMMU, +12.5% on MM-Vet, and +12.9% on OCRBench). We are impressed by the superior results of InternVL2.5 and Qwen2-VL, especially on TextRich. We hypothesize it is due to our i) lower input resolution (e.g., 2.36M vs. 9.63M of InternVL2.5), ii) fewer training stages (e.g., 2 vs. 3 of Qwen2-VL) and iii) frozen vision encoder. This aligns with prior findings (Zhang et al., 2025b) that, when the model size gets larger, higher input resolution and more training stages with fully tunable parameters are pivotal for improving the image performance. Given that our model is video-centric and these enhancements significantly increase training costs, we leave their exploration for future work.
@@ 6 Conclusion
We present SF-LLaVA-1.5, a family of token-efficient Video LLMs for long-form video understanding. Specifically, we integrate the two-stream SlowFast mechanism into a streamlined training pipeline, and perform joint video-image training on a carefully curated data mixture. Our primary focus is on highly efficient model scales (1B and 3B), demonstrating that even relatively small Video LLMs can achieve state-of-the-art performance on video understanding using only publicly available datasets. Experimental results show that SF-LLaVA-1.5 achieves superior results across video benchmarks, while maintaining strong image understanding performance

### v2

@@ Abstract
We introduce SlowFast-LLaVA-1.5 (abbreviated as SF-LLaVA-1.5), a family of video large language models (LLMs) offering a token-efficient solution for long-form video understanding. We incorporate the two-stream SlowFast mechanism into a streamlined training pipeline, and perform joint video-image training on a carefully curated data mixture of only publicly available datasets. Our primary focus is on highly efficient model scales (1B and 3B), demonstrating that even relatively small Video LLMs can achieve state-of-the-art performance on video understanding, meeting the demand for mobile-friendly models. Experimental results demonstrate that SF-LLaVA-1.5 achieves superior performance on a wide range of video and image tasks, with robust results at all model sizes (ranging from 1B to 7B). Notably, SF-LLaVA-1.5 achieves state-of-the-art results in long-form video understanding (e.g., LongVideoBench and MLVU) and excels at small scales across various video benchmarks.
@@ 1 Introduction
Video large language models (LLMs) (Maaz et al., 2024b; Lin et al., 2023a; Xu et al., 2024a) integrate video perception into pre-trained LLMs to process videos and generate responses to user commands. Although significant progress has been made, notable limitations remain in existing Video LLMs. First, they enhance perception and reasoning over long video sequences by leveraging the LLM’s increasing context length and handling massive input frames (Shen et al., 2024; Chen et al., 2024d; Zhang et al., 2024c). However, the potential for transferring this capability to highly efficient models is underexplored. Second, achieving optimal performance typically requires internal datasets and a complex training lifecycle, with selective parameters frozen at each stage (Li et al., 2024a; Zhang et al., 2025a). These intricate designs lead to high computational costs and reproducibility challenges. Third, many Video LLMs (Zohar et al., 2024; Li et al., 2024d) are optimized exclusively for video tasks, limiting their effectiveness as joint models for image understanding tasks.
Building upon the success of SlowFast-LLaVA (Xu et al., 2024b), we introduce SlowFast-LLaVA-1.5, a new family of Video LLMs for long-form video understanding, focusing on the most efficient model scales (1B and 3B). Our model family is both effective and token-efficient in modeling long-range temporal context. This is achieved by employing the SlowFast mechanism, which balances the trade-off between processing more input frames that significantly increases the token count and computational cost, and reducing tokens per frame that inevitably loses fine-grained details. Specifically, the Slow pathway captures detailed spatial features at a low frame rate, while the Fast pathway operates at a high frame rate with fewer tokens per frame to focus on motion cues. The success of our model also relies on a streamlined training pipeline and a carefully curated data mixture. Our model training consists of only two stages. The first stage is supervised fine-tuning on image-only data, providing a good foundation for general knowledge and reasoning. The second stage conducts video-image joint training to learn spatial and temporal features for video understanding while maintaining strong performance in image understanding. To ensure seamless reproducibility, all pre-trained weights and training datasets used in this work are publicly accessible.
@@
Our main contributions are as follows. First, we introduce SlowFast-LLaVA-1.5, a new family of Video LLMs ranging from 1B to 7B parameters. We demonstrate the effectiveness of incorporating the SlowFast mechanism into a supervised fine-tuning framework, modeling long-range context while maintaining high efficiency. Second, our model family provides enhanced reproducibility by using only two training stages and publicly available datasets, distinguishing it from existing methods. Third, SlowFast-LLaVA-1.5 achieves the state-of-the-art performance on long-form video understanding. Moreover, our smaller models (1B and 3B) clearly outperform comparable Video LLMs across video benchmarks.
@@ 2 Related Work
Image Large Language Models have gained widespread attention (Achiam et al., 2023; Team et al., 2023; Touvron et al., 2023; Chen et al., 2024e; Bai et al., 2025). Significant progress across multiple fronts includes: (i) enhancing data quantity and quality during pre-training (McKinzie et al., 2024; Liu et al., 2024a; Lin et al., 2023b; Li et al., 2024c) and supervised fine-tuning (SFT) (Zhang et al., 2025b; Deitke et al., 2024; Chen et al., 2024a; Wang et al., 2023; Tong et al., 2025); (ii) accommodating images of various high resolutions (Lin et al., 2023c; Zhang et al., 2024b; Wang et al., 2024b); (iii) improving architecture designs, including different visual encoders (Zhai et al., 2023; Tong et al., 2024; Shi et al., 2024) and vision-language connectors (Li et al., 2023a; Cha et al., 2024); and iv) conducting comprehensive studies for easy deployment (Team et al., 2023; Marafioti et al., 2025). These rapid advancements also establish a strong foundation for related areas such as video understanding (Maaz et al., 2024b; Lin et al., 2023a), referring & grounding (You et al., 2023; 2024), and visual agents (Durante et al., 2024; Yang et al., 2025).
Video Large Language Models have become an active research area (Li et al., 2023b; Song et al., 2024; Chen et al., 2024b; Zhang et al., 2024e; Zohar et al., 2024). Early Video LLMs are developed as specialist models (Zhang et al., 2023; Cheng et al., 2024; Xu et al., 2024a; Ryoo et al., 2024), achieving strong performance on video tasks but with some trade-offs in image understanding. Training-free Video LLMs (Kim et al., 2024; Xu et al., 2024b) offer an efficient alternative by leveraging Image LLMs without fine-tuning on video data, enabling flexible deployment across various applications. Recent models (Zhang et al., 2024g; Liu et al., 2025; Zhang et al., 2025a) are jointly trained on video and image datasets, obtaining superior results in both modalities. Long-form video understanding (Zhou et al., 2024; Wu et al., 2025) gained increasing attentions, addressing hour-long videos (Chen et al., 2024d; Li et al., 2024d) or live streams (Qian et al., 2024; Zhang et al., 2024a) while optimizing the token efficiency (Lee et al., 2024b). The proposed SlowFast-LLaVA-1.5 is a family of Video LLMs designed for modeling long-range temporal context. It enhances SlowFast-LLaVA (Xu et al., 2024b) by implementing the SlowFast design within a unified video-image training framework, achieving state-of-the-art performance with efficient token utilization.
@@
We provide a detailed explanation of SlowFast-LLaVA-1.5 (abbreviated as SF-LLaVA-1.5), which incorporates the SlowFast video projector into a LLaVA-style architecture, improving long-range temporal modeling while optimizing token efficiency. In contrast to its training-free pioneer (Xu et al., 2024b), this paper (i) systematically investigates different instantiations based on the generic SlowFast idea (Sec. 3.1), (ii) designs a compact yet effective training pipeline (Sec. 3.2), and (iii) introduces tailored data mixtures using only publicly available datasets for each training stage (Sec. 3.3).
@@ 3.1 Model Architecture
As shown in Fig. 1, the architecture of SF-LLaVA-1.5 follows the core design principle of SF-LLaVA (Xu et al., 2024b). It takes a video/image $\mathbf{V}$ and a question $\mathbf{Q}$ as inputs and responds with a textual answer $\mathbf{A}$. For video inputs, we sample $N$ frames, $\mathbf{I}=\{I_{1},I_{2},...,I_{N}\}$, at a fixed frame rate without special frame assembling ($N$ equals $1$ for image input). After that, a visual encoder (e.g., OryxViT (Liu et al., 2025)) is used to extract frame-level features $\mathbf{F}_{v}\in\mathbb{R}^{N\times H\times W}$ from the inputs independently, keeping their original aspect ratio. The video and image feature tokens are then fed into different projectors, with video using the two-stream SlowFast projector and image using a two-layer MLP.
@@ 3.2 Training Pipeline
Stage I (image understanding) conducts SFT with images to provide a good warmup status for video understanding. For simplicity and efficiency, we do not use any extra pretraining stages (Li et al., 2024a) or image splitting strategy (Lin et al., 2023c), although they have proven to be effective for boosting text-rich results. Instead, we use native resolution inputs following Oryx (Liu et al., 2025), where, for each image $I_{i}\in\mathbb{R}^{H_{i}\times W_{i}}$, we have a low resolution $I_{i}^{low}$ and a high resolution $I_{i}^{high}$ input. The low-resolution image is obtained by simply resizing the original image to a base resolution, as in $I_{i}^{low}=\textnormal{resize}(I_{i},H_{i}^{base}\times H_{i}^{base})$. For $I_{i}^{high}$, we keep its original aspect ratio and resize it to $H_{i}^{high}\times W_{i}^{high}$, as in Eq. 1 and 2,
@@ 4.1 Implementation Details
  • Training Stage II continues training based on SF-LLaVA-1.5-Image by combining our video and image data mixture. For image, the high-resolution image is obtained in the same way as Stage I, except that we increase $\theta^{I^{max}}$ to $1536^{2}$. For video, we follow prior work (Zohar et al., 2024) and sample frames at 1 FPS. We set the max frame number to 128 and uniformly sample the frames if the number exceeds this upper bound. For each video frame, we set $\theta^{V^{min}}=288^{2}$ and $\theta^{V^{max}}=480^{2}$. The maximum context length is set to 16K. The models trained by this stage are named as SF-LLaVA-1.5.
@@ 4.3 Image Understanding Results
Second, SF-LLaVA-1.5 outperforms strong baselines at the 7B scale, except for InternVL2.5 and Qwen2-VL. Using MM1.5-7B as an example, SF-LLaVA-1.5 achieves better results across benchmarks (e.g., +7.2% on MMMU, +12.5% on MM-Vet, and +12.9% on OCRBench). We are impressed by the superior results of InternVL2.5 and Qwen2-VL, especially on TextRich. We hypothesize it is due to our (i) lower input resolution (e.g., 2.36M vs. 9.63M of InternVL2.5), (ii) fewer training stages (e.g., 2 vs. 3 of Qwen2-VL) and (iii) frozen vision encoder. This aligns with prior findings (Zhang et al., 2025b) that, when the model size gets larger, higher input resolution and more training stages with fully tunable parameters are pivotal for improving the image performance. Given that our model is video-centric and these enhancements significantly increase training costs, we leave their exploration for future work.
@@ 6 Conclusion
Building upon the insights of SlowFast-LLaVA (Xu et al., 2024b), we introduce SlowFast-LLaVA-1.5, a new family of token-efficient Video LLMs for long-form video understanding. While SlowFast-LLaVA adapts the two-stream SlowFast inputs into a training-free model, this work explores further improvements by building a supervised fine-tuning pipeline with a high-quality data mixture. Our model family, ranging from 1B to 7B parameters, focuses on developing lightweight models that are both compact for potential edge deployment and powerful for various video tasks. Experimental results demonstrate that SlowFast-LLaVA-1.5 achieves superior performance across video benchmarks while maintaining strong image capabilities. We hope our work inspires the community to develop efficient yet robust Long Video LLMs based on open-source datasets.